40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 19 18:10:10 2022, Paco, Update, BHD, Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev DJI_0167.JPGDJI_0169.JPGDJI_0168.JPG
    Reply  Wed Apr 20 08:20:22 2022, Paco, Update, BHD, Part V of BHR upgrade - PR2 weirdness 
       Reply  Wed Apr 20 18:50:03 2022, Koji, Update, BHD, Part V of BHR upgrade - PR2 weirdness Screenshot_2022-04-20_18-46-11.png
    Reply  Thu Apr 21 16:49:01 2022, Paco, Update, BHD, Part V of BHR upgrade - BS Oplev + LO1 offset 
       Reply  Thu Apr 21 18:42:47 2022, Tega, Update, BHD, Part V of BHR upgrade - Align LO and AS beams to BHD BS 
          Reply  Thu Apr 21 20:33:31 2022, Koji, Update, BHD, AS1 UR OSEM problem localized in the chamber Screenshot_from_2022-04-21_20-13-13.png
             Reply  Fri Apr 22 12:03:09 2022, Tega, Update, BHD, AS1 UR OSEM problem localized in the chamber 
                Reply  Fri Apr 22 12:15:08 2022, Anchal, Update, BHD, Cable post installation 
                   Reply  Mon Apr 25 16:57:57 2022, Anchal, Update, BHD, Cable post installation 
Message ID: 16789     Entry time: Wed Apr 20 08:20:22 2022     In reply to: 16788     Reply to this: 16792
Author: Paco 
Type: Update 
Category: BHD 
Subject: Part V of BHR upgrade - PR2 weirdness 

Yesterday, I tried tuning the PR2 damping gains by increasing the gain until the damping gave the ~5 oscillations (by watching the damped motion using StripTool, and keeping an eye on the PD var). I noticed that often when I changed the gain, some OSEM sensors shifted (gained an offset!!) and the PD var values changed, typically increasing at higher damping gains. I reverted the changes until the PD var looked "normal" again (~ 2 mV) but it is hard to imagine that the damping filters can have such a "DC" effect, given the shape comprises single zero at 0 Hz (and pole at 30 Hz).

ELOG V3.1.3-